OpenStudy (anonymous):

need to find the integral of the differention so -integral dv from 0 to 10 not sure what to do with the dv

OpenStudy (unklerhaukus):

\[-\int\limits_0^{10}\text dv \]?

OpenStudy (unklerhaukus):

it this what you are trying to do./?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

ok so the trick is that \(\text dv\) can be written as \(1\cdot\text dv\) ..... as \(v^0\cdot\text dv\)

OpenStudy (unklerhaukus):

so when we integrate, we add one to the index, and divide by the new index

OpenStudy (anonymous):

What's the index

OpenStudy (unklerhaukus):

the power or index of \(x^n\) is the n bit

OpenStudy (unklerhaukus):

or exponent

OpenStudy (unklerhaukus):

in this case it is zero

OpenStudy (unklerhaukus):

\[-\int\limits_0^{10}\text dv=-\int\limits_0^{10}v^0\cdot\text dv=-\left.\frac {v^{0+1}}{0+1}\right|_0^{10}=\]